tne end OT ZU18, witn a Tew pieces missin Not So Strong, Inc. Income Statement Service Revenues $56,350 Expenses Utilities Wages Rent $800 $22,100 $18,000 NET INCOME Not So Strong, Inc. Statement of Stockholders Equity Total Retained Stockholders Equity $17,720 $6,000 Common Stock (CS) Earnings $13,720 Beginning Balance CS Issuances $6,000 Add: Net Income Less: Dividends Ending Balance $36,170 Not So Strong, Inc. Balance Sheet Assets: Cash A/R Supplies $23,490 TOTAL ASSETS: Liabilities: Accounts Payable Wages Payable Total Liabilities Stockholders Equity: Common Stock Retained Earnings $36,170 $61,150 Total Stockholders Equity TOTAL LIABILITIES AND STOCKHOLDERS EQUITY: Calculate the missing amounts
Answer #1

Answer:

Not so Strong, Inc.
Income Statement
Service Revenues                               56,350
Expenses Utilities
Utilities                                     800
Wages                               22,100
Rent                               18,000
NET INCOME(A)($56,350-$800-$22,100-$18,000)                               15,450
Not So Strong, Inc.
Statement of Stockholders' Equity
Common Stock (CS) Retained Earnings Total Retained Stockholders' Equity
Beginning Balance C=$4,000               13,720                              17,720
CS Issuances                                  6,000                        -                                  6,000
Add: Net Income               15,450                              15,450
Less: Dividends B=$3,000 B=$3,000
Ending Balance D=$10,000 E=$26,170                              36,170
Not So Strong, Inc.
Balance Sheet
Assets Amt($)
Cash f=35,080
A/R                               23,490
Supplies                                  2,580
Total Assets: g=61,150
Liabilities:
Account Payable h=22,660
Wages Payable                                  2,320
Total Liabilities i=24,900
Stockholders' Equity:
Common Stock D=$10,000
Retained Earnings E=$26,170
Total Stockholders' Equity                               36,170
TOTAL LIABILITIES AND STOCKHOLDERS' EQUITY                               61,150
Answer: Amt($)
(A)=($56,350-$800-$22,100-$18,000)                               15,450
B=(36,170-15,450-6,000-17,720)                                  3,000
C = 17,720-13,720                                  4,000
D= 4,000+6,000                               10,000
E= 36,170-10,000                               26,170
F= 61,170-23,490-2,580                               35,080
G=Total Assets=TOTAL LIABILITIES AND STOCKHOLDERS' EQUITY                               61,150
H=24,900-2,320                               22,660
I= 61,150-36,150                               24,900